# A Good Foundation is Worth Many Labels: Label-Efficient Panoptic Segmentation

### v1

@@ IV-C Ablations and Analysis
Choice of Labeled Images. To measure the effect of the selected images, we conduct a user study with four participants tasked to select ten RGB images covering all semantic classes while maximizing diversity. We report the mean and standard deviation in Tab. VI showing that the performance of PASTEL remains stable for different selections of training data. Next, we evaluate the potential performance of PASTEL if one would further increase the number of labeled training images. Please note that we use the same images as selected by Käppeler et al. [13] allowing for direct comparison with their study. Without adapting any further parameters, PASTEL achieves a mIoU of and a PQ of when using 100 images, i.e., an improvement of mIoU to the previous state of the art listed in Tab. I.
Backbone. We present results for different variants of DINOv2 [15] in Tab. VII. Note that we selected DINOv2 ViT-B/14 for our method as it compromises performance and computational feasibility. We observe that the larger backbone DINOv2 ViT-L/14 shows significant performance improvements, whereas increases due to DINOv2 ViT-g/14 are marginal. This aligns with prior studies on image classification by the authors of DINOv2 [15].
Iterative Self-Training. Finally, we conduct ablation studies on the number of self-training iterations (Tab. VIII) and images sampled by our sampling strategy (Tab. IX). We confirm previous findings [36] and employ only a single iteration of self-training. Concerning the image count, we sample the nearest neighbors for each annotated image in the training set and continue training the semantic head for 50 epochs. In Tab. IX, we further report results for , where corresponds to resuming the training without pseudo-labeled images. Note that our proposed similarity sampling yields significantly better results than self-training on randomly sampled images.

### v2

@@ IV-C Ablations and Analysis
Each row also includes the components of all rows above.
@@
Choice of Labeled Images. To measure the effect of the selected images, we conduct a user study with four participants tasked to select ten RGB images covering all semantic classes while maximizing diversity. We report the mean and standard deviation in the bottom row of Tab. VI denoted by study. The study shows that the performance of PASTEL remains stable for different selections of training data. Next, we evaluate the potential performance of PASTEL if one would further increase the number of labeled training images, although this does not reflect the main goal of our work. Please note that we use the same images as SPINO [13], allowing for a direct comparison. In detail, PASTEL achieves , , , , and PQ compared to SPINO for an increasing label count from to .
Backbone. We present results for different variants of DINOv2 [15] in Tab. VII. Note that we selected DINOv2 ViT-B/14 for our method as it compromises performance and computational feasibility. We observe that the larger backbone DINOv2 ViT-L/14 shows significant performance improvements, whereas increases due to DINOv2 ViT-g/14 are marginal. Similar to a study on image classification [15], we hypothesize that the number of parameters of the ViT-L/14 variant suffices to model the training data.
Iterative Self-Training. Finally, we conduct studies on the number of self-training iterations (Tab. VIII) and images sampled by our sampling strategy (Tab. IX). We show that performing self-training once is sufficient to increase performance [37], while further iterations decrease the quality, most likely due to overfitting. For the image count, we sample the nearest neighbors for each annotated image in the training set and continue training the semantic head for 50 epochs. In Tab. IX, we further report results for , where corresponds to resuming the training without pseudo-labeled images. Note that our proposed similarity sampling yields significantly better results than self-training on randomly sampled images, shown in the bottom row.